# Silence Speaks Volumes: A New Paradigm for Covert Communication via History Timing Patterns

### v1

@@ 1 Introduction
The landscape of network cybersecurity is constantly evolving, marked by the perpetual back and forth between attackers seeking covert communication and defenders striving to lock down network environments [NIHBook]. A persistent challenge in cybersecurity is the utilization and detection of Covert Channel (CC)s, which circumvent security policies by embedding hidden communications within legitimate network operations [petitcolas1999information, Schmid23, zander2007survey]. CC may be used for malware command and control or data exfiltration [NIHBook, caviglione2022never, zander2007survey], but also by journalists looking to report under surveillance [wustrow2011telex, fifield2012evading] and citizens aiming to bypass Internet censorship [AsiaCCS24:OPPR]. The primary impact of a CC has been traditionally measured by the channel’s number of bits transferred per second (bandwidth) [twenty02].
CC can be understood as steganographic methods that modify overt information objects using different hiding patterns to hide secret information from plain view [NIHBook, zander2007survey]. Traditional covert channels often rely on the modification of network traffic patterns or the timing of packet transmissions to embed hidden messages [mileva14, wendzelpatternsurvey15]. However, these methods have been regularly investigated during the last 20 years and are now increasingly susceptible to detection by traffic analysis tools and intrusion detection systems, using methods such as anomaly detection [detect22].
A recent advancement in this domain is the introduction of the History Covert Channel (HCC[DYST22]. They minimize the detectable footprint of covert communication by referencing legitimate network traffic through small pointers rather than creating or modifying cover traffic. This approach enables an amplification of the covert message to be transferred (transferring a small pointer instead of the larger bitstring being pointed to). One such protocol suggested by Wendzel et al. is Did You See That (DYST) [DYST22]. While it is protocol-agnostic in principle, its implementation leverages ARP requests to reference hashes of overt network traffic. Despite its innovative approach, DYST is not without limitations, particularly in terms of detection risk due to signaling based on overt traffic, its dependence on clock synchronization, and scalability in high packet rate environments.
@@ 2 Background and Related Work
CCs are hidden communication paths that allow the transfer of information in a manner that circumvents system security policies [wendzelpatternsurvey15]. Traditionally, network-based covert channels have embedded secret messages in network communications by manipulating packet sizes, timing, or other traffic characteristics. Since the introduction of CC by Lampson in 1973 [lamp73], researchers have focused on two primary methods [DYST22]: (1) active sending, where the covert sender creates its own traffic to embed secret data, and (2) passive sending, where legitimate traffic is modified to carry covert information. Both methods have a common limitation — they render the covert communication detectable due to the unnatural patterns they introduce into the network traffic.
@@
Traditional network CCs embed information within the timing, packet sequence, or other characteristics of transmissions, and several methods are difficult to detect using conventional security measures [mileva14, wendzelpatternsurvey15, Iglesias2017Are, wang2016dwt, NIHBook, epskappa24].
@@
Network CCs based on encoding data in timing information include Inter-Packet Delay (IPD)-based ones such as Jitterbug [jitterbug16], which use the time difference between packets as input, and packet-sorting-based encoding, which utilizes the order of a particular set of packets as a channel [wendzelpatternsurvey15, cabuk04]. These approaches can also be combined into more complex encoding methods.
@@
All of them may use the timing information of all packets from a given connection, or focus on specific Packets of Interest (POI), such as Address Resolution Protocol (ARP) requests, DNS queries, or other common protocol messages [wendzelpatternsurvey15]. POIs may be chosen based on their attributes–such as length, timing, or classification–to optimize throughput, preserve legitimate network functionality, or avoid detection.
@@
Several approaches have aimed to reduce attention to covert transmissions [NIHBook, wendzelLowAttention]. A recent development led to CCs that are capable of transmitting high-bandwidth data over encrypted channels by taking over and controlling encrypted overt communication [cckex24]. But in some cases even a CC in encrypted traffic can be detected using machine learning on statistics such as the source address and the entropy of the timestamp [ai24, morepackets21, Ondov2022Covert, Epishkina2019Timing]. The likelihood of being able to discern overt and covert channels using such statistics increases as the ratio between covert and overt signals increases. This holds even for methods that specifically focus on detection in small samples [GAS22].
@@
History CCs represent an advancement in this regard, because they communicate almost exclusively by pointing to existing legitimate network traffic rather than altering or generating new traffic. This allows them to amplify the number of bits transferred in the channel by sending only a pointer covertly, as can be seen in Fig. 1. Amplification minimizes observable artifacts, making HCC a promising avenue for evading modern detection systems while keeping the net amount of data transferred the same [DYST22, wang5634454moss]. The ratio of bits effectively received compared to the number of bits used for the pointer is called the Covert Amplification Factor (CAF[DYST22]:
@@
The pointer refers the receiver to unaltered legitimate traffic, whose hash value is then used as bits of the secret message. DYST uses Protocol Data Unit (PDU) datafields like IP checksum as the input value in combination with the PDU’s timestamp. This enables DYST to use the potentially high entropy of timestamps for a data input source. However, this potential can only be realized in practice if the involved systems can reliably synchronize their clocks with a very high degree of accuracy – ideally down to the same millisecond. DYST relies on other protocols such as NTP for clock synchronization [DYST22]. Since prior studies suggest that NTP synchronization of internet systems can typically vary between tens of milliseconds [ntp13] and a few milliseconds [ntp19], this need for clock synchronization limits DYST’s effectiveness, even if the covert agents themselves are close to each other, as is visualized in Fig. 2. If the covert agents are not overt systems themselves, their NTP synchronization would also be a suspicious signal in itself. Furthermore, clock drift or NTP errors can lead to desynchronization, compromising the effectiveness and reliability of the channel, which are essential for covert communication.
@@ 3.1.1 Adversary Capabilities
SHP operates within the traditional threat model for a covert network channel, where an adversary monitors network traffic but does not have access to endpoints [Vivek2014Improving, Liu2013RealTime]. The adversary is assumed to have extensive network monitoring capabilities and actively searches for CC by analyzing the characteristics of network traffic [GAS22, Ayub2019A]. In particular, the adversary aims to detect covert communication based on timing patterns by employing statistical detection thresholds, such as identifying timing deviations or unusual traffic volumes from certain systems [Ondov2022Covert, Iglesias2017Are].
@@
Traditional detection approaches for timing-based covert channels focus on identifying statistical anomalies in packet inter-arrival times or increased transmission delays caused by embedding hidden information [Iglesias2017Are, Chen2015A, Djidjev2011Graph], often combined with machine learning techniques [Darwish2019Using, Epishkina2019Timing]. Prior research suggests that a basic timing-based CC can be detected if its manipulations result in an increase of the average delay by a factor of two or more [aleidi20]. Recent advances have produced more sophisticated covert channels that evade detection. For example, the Epsilon-Kappa-libur[epskappa24] introduces an adaptive encoding mechanism that circumvents popular heuristics and degrades the performance of recent ML-based detection methods such as GAS and SnapCatch[GAS22, snapcatch20]. In response, detection techniques have also evolved, with recent work using isolated binary trees advancing the detection of such sophisticated covert channels [LIN2025104200].
@@ 3.1.2 Defense Strategy
To achieve this, we leverage the principle of channel amplification first developed in DYST [DYST22]. Amplification minimizes the footprint of covert signaling by encoding information via small pointers to legitimate network events rather than modifying the timing or content of packets directly. This design philosophy ensures that even sophisticated statistical analyses would struggle to differentiate SHP signals from normal network variations, as the actual modifications to traffic patterns are minimal compared to the volume of information transferred. Furthermore, by carefully calibrating the amplification factor, we can maintain a favorable trade-off between covert bandwidth and detectability risk.
@@ 3.1.3 Operational Scenarios
We consider two scenarios that typify the application areas of a covert network channel [Schneider2023Network, Heinz2020Covert]: LAN and WAN.
@@
In the LAN scenario, covert agents filter overt traffic to use only packets that are visible to all network nodes as POI, ensuring they can be observed by both the CS and CR when residing in the same broadcast domain network [Schmid23, dua2021covert]. The proximity of communicating nodes in this scenario offers lower latency and more predictable network conditions, but also poses different challenges as broadcast traffic volumes may be lower and more scrutinized by local security monitoring. Furthermore, local network segments often implement more stringent traffic filtering and analysis, requiring SHP to adapt its parameters to maintain stealthiness while achieving acceptable bandwidth.
@@ 3.2.1 Basic algorithm
SHP uses timing information that is observable by all covertly communicating nodes as the input source for a history protocol. To improve the probability of matching a random bit distribution, the input value is deskewed before matching, e.g. by hashing it [cachin98].
@@
The expected number of attempts to hit the chosen sequence follows a geometric distribution [beria2006confidence]. The expected value of a geometric distribution with success probability $p$ is $1/p$. Therefore, the expected number of attempts $EA$ to hit the chosen sequence is:
@@ 3.2.2 Packets of Interest (POI)
In the LAN scenario, only overt traffic that can reasonably be expected to be visible to all network nodes is used as POI. For this, DYST relied on Ethernet and Internet Protocol (IP) broadcast traffic, as well as IPv6 link-local multicast (e.g., ff02::/16[DYST22].
@@ 3.2.3 Network-timing input sources
In contrast to CC based on actively manipulating time pattern, our algorithm will have to rely on a source of input bits that is already available on the network. It follows that, in order to optimize the data carrying capacity of the channel, it is essential to use a readily available input source with substantial informational density. In other words, the input source should have a preexisting high entropy as established by Shannon [shannon48], since it quantifies the uncertainty or the average informativeness of a variable’s possible outcomes. Therefore, entropy, given by
@@
Although clock timestamps have the necessary entropy, nearly identical timestamps between sender and receiver are an additional requirement for SHP to function properly, because this assumes fine-grained clock synchronization between communicating nodes and limits communication to connections that can ensure minimal delays during transmission. This was also a limiting factor for DYST, especially in the WAN scenario, where agents may not even observe POI at roughly the same time, much less send and receive covert data within tight timing windows. Conceptually, embracing uncertainty in the observable timing, rather than fighting it via tight wall-clock synchronization, echoes recent covert designs that randomize transmission instants to maintain covertness under a warden’s binary test[Lu2023TimeUncertainty].
@@
This approach simplifies synchronization by eliminating the need for continuous external time corrections, thus preserving the high entropy of the timing data while enhancing the overall robustness and applicability of SHP across various network environments. This is especially true if timestamps need to be rounded, since this can easily nullify any IPDs in congested environments. Because an initial synchronization signal is needed to initialize a connection regardless [cabuk06], this approach is also not more detectable than using other sources of input data.
One caveat to using the relative timing approach in the LAN scenario is that more POI might arrive at the CR before the pointer has been calculated and transmitted. This would lead to the CR interpreting the pointer as pointing to the wrong packet. Such a POI mismatch is less problematic for SHP than it was for DYST, because we do not use the POIs data bits for interpretation, but only their timing. As long as the measured timing differences between CS and CR stay within the allotted rounding window, the interpretation of the message will still be correct. To mitigate this issue further, several approaches can be used. First, more aggressive rounding leads to more tolerant matching. Second, instead of a naive implementation that uses the last POI that arrived at the CR, the pointer delay can be retroactively subtracted from the arrival time to guess the right POI. This requires determining the round trip time between agents at the start of the connection and then subtracting $\frac{rtt}{2}$ at the CR from pointer arrival times. Third, agents can ensure a sufficient IPD between POI by enforcing a silence interval, a method recommended by Wendzel et al. [DYST22]. Because POI often occur in bursts, silence intervals lasting more than a few milliseconds substantially reduce the number of POI used substantially. Lastly, we can actually remove the need to reference the right POI entirely by directly using the IPD between both pointers. Even when using this direct reference the matching on the CS only happens when matching POI are detected, so the timing characteristics of the covert channel stay anchored to overt traffic, ensuring undetectability stays stable.
@@ 3.2.4 Algorithm Extensions
Modern wired networks generally exhibit low bit error rates; however, wireless networks face higher susceptibility due to signal attenuation, interference, and multipath propagation [jacobs2015]. Given the additional noise inherent in network timing, robust error correction mechanisms are essential for reliable transmission.
@@
Rehashing is a feature for HCC first introduced in DYST [DYST22]. Without this feature, a single hash is computed per POI. Throughput can be enhanced by introducing multiple hashing attempts, allowing a secret message to be discovered by computing multiple hashes up to a given rehashing budget of bits $m$. This permits up to $2^{m}$ of additional hashing attempts per POI, icreasing the likehood of a match per POI. This approach requires us to inform the receiver exactly how many hashing iterations are needed to find the data pointed to. Rehashing therefore also reduces the CAF or requires increasing the bitlengthbitlength parameter to keep it stable.
@@ 3.3 Choosing parameters
A CC can be characterized by its position in the magic triangle of the three basic goals covert bandwidth, robustness and undetectability [fridrich1999applications]. The selection of optimal parameters depends on the specific constraints and goals of the covert communication scenario. If undetectability is paramount, the CAF may be prioritized by increasing bitlengthbitlength, which will in turn decrease the number of pointer signals required to transmit a given message.

### v2

@@ Abstract
This is a pre-print.
The final and enhanced version of this paper was published under a different title in the Journal of Information Security and Applications (JISA): https://doi.org/10.1016/j.jisa.2026.104431
@@ 1 Introduction
The landscape of network cybersecurity is constantly evolving, marked by the perpetual back and forth between attackers seeking covert communication and defenders striving to lock down network environments [43]. A persistent challenge in cybersecurity is the utilization and detection of Covert Channel (CC)s, which circumvent security policies by embedding hidden communications within legitimate network operations [48, 49, 68]. CC may be used for malware command and control or data exfiltration [43, 12, 68], but also by journalists looking to report under surveillance [64, 19] and citizens aiming to bypass Internet censorship [72]. The primary impact of a CC has been traditionally measured by the channel’s number of bits transferred per second (bandwidth) [32].
CC can be understood as steganographic methods that modify overt information objects using different hiding patterns to hide secret information from plain view [43, 68]. Traditional covert channels often rely on the modification of network traffic patterns or the timing of packet transmissions to embed hidden messages [44, 63]. However, these methods have been regularly investigated during the last 20 years and are now increasingly susceptible to detection by traffic analysis tools and intrusion detection systems, using methods such as anomaly detection [24].
A recent advancement in this domain is the introduction of the History Covert Channel (HCC[62]. They minimize the detectable footprint of covert communication by referencing legitimate network traffic through small pointers rather than creating or modifying cover traffic. This approach enables an amplification of the covert message to be transferred (transferring a small pointer instead of the larger bitstring being pointed to). One such protocol suggested by Wendzel et al. is Did You See That (DYST) [62]. While it is protocol-agnostic in principle, its implementation leverages ARP requests to reference hashes of overt network traffic. Despite its innovative approach, DYST is not without limitations, particularly in terms of detection risk due to signaling based on overt traffic, its dependence on clock synchronization, and scalability in high packet rate environments.
@@ 2 Background and Related Work
CCs are hidden communication paths that allow the transfer of information in a manner that circumvents system security policies [63]. Traditionally, network-based covert channels have embedded secret messages in network communications by manipulating packet sizes, timing, or other traffic characteristics. Since the introduction of CC by Lampson in 1973 [34], researchers have focused on two primary methods [62]: (1) active sending, where the covert sender creates its own traffic to embed secret data, and (2) passive sending, where legitimate traffic is modified to carry covert information. Both methods have a common limitation — they render the covert communication detectable due to the unnatural patterns they introduce into the network traffic.
@@
Traditional network CCs embed information within the timing, packet sequence, or other characteristics of transmissions, and several methods are difficult to detect using conventional security measures [44, 63, 27, 59, 43, 73].
@@
Network CCs based on encoding data in timing information include Inter-Packet Delay (IPD)-based ones such as Jitterbug [58], which use the time difference between packets as input, and packet-sorting-based encoding, which utilizes the order of a particular set of packets as a channel [63, 8]. These approaches can also be combined into more complex encoding methods.
@@
All of them may use the timing information of all packets from a given connection, or focus on specific Packets of Interest (POI), such as Address Resolution Protocol (ARP) requests, DNS queries, or other common protocol messages [63]. POIs may be chosen based on their attributes–such as length, timing, or classification–to optimize throughput, preserve legitimate network functionality, or avoid detection.
@@
Several approaches have aimed to reduce attention to covert transmissions [43, 61]. A recent development led to CCs that are capable of transmitting high-bandwidth data over encrypted channels by taking over and controlling encrypted overt communication [38]. But in some cases even a CC in encrypted traffic can be detected using machine learning on statistics such as the source address and the entropy of the timestamp [29, 46, 47, 18]. The likelihood of being able to discern overt and covert channels using such statistics increases as the ratio between covert and overt signals increases. This holds even for methods that specifically focus on detection in small samples [35].
@@
History CCs represent an advancement in this regard, because they communicate almost exclusively by pointing to existing legitimate network traffic rather than altering or generating new traffic. This allows them to amplify the number of bits transferred in the channel by sending only a pointer covertly, as can be seen in Fig. 1. Amplification minimizes observable artifacts, making HCC a promising avenue for evading modern detection systems while keeping the net amount of data transferred the same [62, 60]. The ratio of bits effectively received compared to the number of bits used for the pointer is called the Covert Amplification Factor (CAF[62]:
@@
The pointer refers the receiver to unaltered legitimate traffic, whose hash value is then used as bits of the secret message. DYST uses Protocol Data Unit (PDU) datafields like IP checksum as the input value in combination with the PDU’s timestamp. This enables DYST to use the potentially high entropy of timestamps for a data input source. However, this potential can only be realized in practice if the involved systems can reliably synchronize their clocks with a very high degree of accuracy – ideally down to the same millisecond. DYST relies on other protocols such as NTP for clock synchronization [62]. Since prior studies suggest that NTP synchronization of internet systems can typically vary between tens of milliseconds [21] and a few milliseconds [45], this need for clock synchronization limits DYST’s effectiveness, even if the covert agents themselves are close to each other, as is visualized in Fig. 2. If the covert agents are not overt systems themselves, their NTP synchronization would also be a suspicious signal in itself. Furthermore, clock drift or NTP errors can lead to desynchronization, compromising the effectiveness and reliability of the channel, which are essential for covert communication.
@@ 3.1.1 Adversary Capabilities
SHP operates within the traditional threat model for a covert network channel, where an adversary monitors network traffic but does not have access to endpoints [57, 39]. The adversary is assumed to have extensive network monitoring capabilities and actively searches for CC by analyzing the characteristics of network traffic [35, 5]. In particular, the adversary aims to detect covert communication based on timing patterns by employing statistical detection thresholds, such as identifying timing deviations or unusual traffic volumes from certain systems [47, 27].
@@
Traditional detection approaches for timing-based covert channels focus on identifying statistical anomalies in packet inter-arrival times or increased transmission delays caused by embedding hidden information [27, 13, 16], often combined with machine learning techniques [15, 18]. Prior research suggests that a basic timing-based CC can be detected if its manipulations result in an increase of the average delay by a factor of two or more [2]. Recent advances have produced more sophisticated covert channels that evade detection. For example, the Epsilon-Kappa-libur[73] introduces an adaptive encoding mechanism that circumvents popular heuristics and degrades the performance of recent ML-based detection methods such as GAS and SnapCatch[35, 1]. In response, detection techniques have also evolved, with recent work using isolated binary trees advancing the detection of such sophisticated covert channels [37].
@@ 3.1.2 Defense Strategy
To achieve this, we leverage the principle of channel amplification first developed in DYST [62]. Amplification minimizes the footprint of covert signaling by encoding information via small pointers to legitimate network events rather than modifying the timing or content of packets directly. This design philosophy ensures that even sophisticated statistical analyses would struggle to differentiate SHP signals from normal network variations, as the actual modifications to traffic patterns are minimal compared to the volume of information transferred. Furthermore, by carefully calibrating the amplification factor, we can maintain a favorable trade-off between covert bandwidth and detectability risk.
@@ 3.1.3 Operational Scenarios
We consider two scenarios that typify the application areas of a covert network channel [50, 25]: LAN and WAN.
@@
In the LAN scenario, covert agents filter overt traffic to use only packets that are visible to all network nodes as POI, ensuring they can be observed by both the CS and CR when residing in the same broadcast domain network [49, 17]. The proximity of communicating nodes in this scenario offers lower latency and more predictable network conditions, but also poses different challenges as broadcast traffic volumes may be lower and more scrutinized by local security monitoring. Furthermore, local network segments often implement more stringent traffic filtering and analysis, requiring SHP to adapt its parameters to maintain stealthiness while achieving acceptable bandwidth.
@@ 3.2.1 Basic algorithm
SHP uses timing information that is observable by all covertly communicating nodes as the input source for a history protocol. To improve the probability of matching a random bit distribution, the input value is deskewed before matching, e.g. by hashing it [11].
@@
The expected number of attempts to hit the chosen sequence follows a geometric distribution [6]. The expected value of a geometric distribution with success probability $p$ is $1/p$. Therefore, the expected number of attempts $EA$ to hit the chosen sequence is:
@@ 3.2.2 Packets of Interest (POI)
In the LAN scenario, only overt traffic that can reasonably be expected to be visible to all network nodes is used as POI. For this, DYST relied on Ethernet and Internet Protocol (IP) broadcast traffic, as well as IPv6 link-local multicast (e.g., ff02::/16[62].
@@ 3.2.3 Network-timing input sources
In contrast to CC based on actively manipulating time pattern, our algorithm will have to rely on a source of input bits that is already available on the network. It follows that, in order to optimize the data carrying capacity of the channel, it is essential to use a readily available input source with substantial informational density. In other words, the input source should have a preexisting high entropy as established by Shannon [51], since it quantifies the uncertainty or the average informativeness of a variable’s possible outcomes. Therefore, entropy, given by
@@
Although clock timestamps have the necessary entropy, nearly identical timestamps between sender and receiver are an additional requirement for SHP to function properly, because this assumes fine-grained clock synchronization between communicating nodes and limits communication to connections that can ensure minimal delays during transmission. This was also a limiting factor for DYST, especially in the WAN scenario, where agents may not even observe POI at roughly the same time, much less send and receive covert data within tight timing windows. Conceptually, embracing uncertainty in the observable timing, rather than fighting it via tight wall-clock synchronization, echoes recent covert designs that randomize transmission instants to maintain covertness under a warden’s binary test[41].
@@
This approach simplifies synchronization by eliminating the need for continuous external time corrections, thus preserving the high entropy of the timing data while enhancing the overall robustness and applicability of SHP across various network environments. This is especially true if timestamps need to be rounded, since this can easily nullify any IPDs in congested environments. Because an initial synchronization signal is needed to initialize a connection regardless [10], this approach is also not more detectable than using other sources of input data.
One caveat to using the relative timing approach in the LAN scenario is that more POI might arrive at the CR before the pointer has been calculated and transmitted. This would lead to the CR interpreting the pointer as pointing to the wrong packet. Such a POI mismatch is less problematic for SHP than it was for DYST, because we do not use the POIs data bits for interpretation, but only their timing. As long as the measured timing differences between CS and CR stay within the allotted rounding window, the interpretation of the message will still be correct. To mitigate this issue further, several approaches can be used. First, more aggressive rounding leads to more tolerant matching. Second, instead of a naive implementation that uses the last POI that arrived at the CR, the pointer delay can be retroactively subtracted from the arrival time to guess the right POI. This requires determining the round trip time between agents at the start of the connection and then subtracting $\frac{rtt}{2}$ at the CR from pointer arrival times. Third, agents can ensure a sufficient IPD between POI by enforcing a silence interval, a method recommended by Wendzel et al. [62]. Because POI often occur in bursts, silence intervals lasting more than a few milliseconds substantially reduce the number of POI used substantially. Lastly, we can actually remove the need to reference the right POI entirely by directly using the IPD between both pointers. Even when using this direct reference the matching on the CS only happens when matching POI are detected, so the timing characteristics of the covert channel stay anchored to overt traffic, ensuring undetectability stays stable.
@@ 3.2.4 Algorithm Extensions
Modern wired networks generally exhibit low bit error rates; however, wireless networks face higher susceptibility due to signal attenuation, interference, and multipath propagation [28]. Given the additional noise inherent in network timing, robust error correction mechanisms are essential for reliable transmission.
@@
Rehashing is a feature for HCC first introduced in DYST [62]. Without this feature, a single hash is computed per POI. Throughput can be enhanced by introducing multiple hashing attempts, allowing a secret message to be discovered by computing multiple hashes up to a given rehashing budget of bits $m$. This permits up to $2^{m}$ of additional hashing attempts per POI, icreasing the likehood of a match per POI. This approach requires us to inform the receiver exactly how many hashing iterations are needed to find the data pointed to. Rehashing therefore also reduces the CAF or requires increasing the bitlengthbitlength parameter to keep it stable.
@@ 3.3 Choosing parameters
A CC can be characterized by its position in the magic triangle of the three basic goals covert bandwidth, robustness and undetectability [20]. The selection of optimal parameters depends on the specific constraints and goals of the covert communication scenario. If undetectability is paramount, the CAF may be prioritized by increasing bitlengthbitlength, which will in turn decrease the number of pointer signals required to transmit a given message.